Concord Pet Care Clinic paid $150,000 for a group purchase of land, building, and equipment. At the time of the acquisition, the land had a market value of $80,000, the building $64,000, and the equipment $16,000. Journalize the lump-sum purchase of the three assets for a total cost of $150,000, the amount for which the business signed a note payable. (Record a single compound journal entry. Record debits first, then credits. Select the explanation on the last line of the journal entry table.) Date Accounts and Explanation Debit Credit
Answer #1

Ans:

Journal Entry

The journal entry to record the lump-sum purchase:

Date

Accounting Entry and Explanation

Debit

Credit

Land

$75000

Building

$60000

Equipment

$15000

                Cash

$150000

Calculations:

Particulars

Amount

Land

$80,000

Building

$64,000

Equipment

$16,000

Total Fair Value

$160,000

Formula

Fixed assets purchased for lump-sum are therefore recorded at a value calculated using their respective market values using total market value

Value of Asset =

                 Fair value of each asset

× Lump sum Amount Paid

Total Fair value of assets

Value calculation:

Land=150,000*80,000/160,000=$75,000

Building=$150,000*64,000/160,000=$60,000

Equipment=$150,000*16000/160,000=$15000